# Dereflection Any Image with Diffusion Priors and Diversified Data

### v1

@@ Appendix A Details of Loss Functions
The noise introduced during the multi-step diffusion process creates challenges in producing stable results, making traditional diffusion models unsuitable for our reflection removal task, which requires deterministic outputs. In contrast, our approach employs one-step denoising [49, 46], which significantly accelerates inference speed while producing deterministic results. The denoising loss function for our method is formulated as:
@@
We adapt ControlNet [50] as a mechanism to inject information from mixed images. By jointly training the ControlNet with the upsampling blocks of the U-Net architecture, we enable the model to effectively perform reflection removal. Our loss function is formulated as:
@@
At the final stage of our progressive training, the cross-latent decoder is trained using a combination of L1 loss, SSIM [39] loss, and LPIPS [51] loss, with equal weighting applied to each component:
@@
Appendix B More Qualitative Comparison
We provide additional qualitative comparisons between our method and other approaches, as illustrated in Fig. 10 and Fig. 11. Notably, the other approaches are not finetuned on our data setting, which highlights the superior performance of our comprehensive method.
Appendix C Additional examples of DRR
To show the diversity of our dataset, we randomly select several mixed images of one scene, demonstrated in Fig. 12. The ground truth transmission layer is the first image while the remaining images represent various mixed images.
Appendix D Downstream application
@@
Figure 10: Qualitative comparison of our method with other approaches (not finetuned on our data setting). The benchmark datasets, listed from top to bottom, are $Nature$, $Real$, $SIR^{2}$.
Figure 11: Qualitative comparison of our method with other approaches (not finetuned on our data setting). The benchmark datasets, listed from top to bottom, are $DRR$-$S$ and $DRR$-$C$.

### v2

@@ Appendix A Details of Loss Functions
The noise introduced during the multi-step diffusion process creates challenges in producing stable results, making traditional diffusion models unsuitable for our reflection removal task, which requires deterministic outputs. In contrast, our approach employs one-step denoising (Ye et al. 2024; Xu et al. 2024), which significantly accelerates inference speed while producing deterministic results. The denoising loss function for our method is formulated as:
@@
We adapt ControlNet (Zhang, Rao, and Agrawala 2023) as a mechanism to inject information from mixed images. By jointly training the ControlNet with the upsampling blocks of the U-Net architecture, we enable the model to effectively perform reflection removal. Our loss function is formulated as:
@@
At the final stage of our progressive training, the cross-latent decoder is trained using a combination of L1 loss, SSIM (Wang et al. 2004) loss, and LPIPS (Zhang et al. 2018) loss, with equal weighting applied to each component:
@@
Appendix B More Details of DRR
To better illustrate the diversity of our dataset, we have included a pie chart (Fig. 9) visualizing the distribution of scene types, ranging from indoor scenes, mordern and natural outdoor scenes and object-centric scenes. This highlights the balanced and comprehensive coverage of our data. To show the angle diversity of our dataset, we randomly select several mixed images of one scene, demonstrated in Fig. 15. The ground truth transmission layer is the first image while the remaining images represent various mixed images.
Figure 9: Scene types of DRR, highlighting the balanced and comprehensive coverage of our data.
Appendix C More Details of Synthetic Data
In addition to datasets collected from the real world, synthetic data can serve as a significant supplementary resource to enrich the diversity of training data. In our work, we adopt the following formulation from DSRNet (Hu and Guo 2023) to generate synthetic data:
where T, R, I represent the transmission, reflection and mixed layers, respectively. To enhance the diversity of the synthesized data, we randomly sample transmission and reflection layers from the COCO (Lin et al. 2014) and PASCAL VOC (Everingham et al. 2015) datasets and assign random values to the coefficients $\gamma_{1}\in[0.8,1.0]$ and $\gamma_{2}\in[0.4,1.0]$ during synthesis. An intuitive observation is that the synthetic data exhibit a wide range of quality: while some images closely resemble real-world reflections, others appear less realistic. To address this, we leverage CLIP similarity (Radford et al. 2021) with the text prompt “image with glass reflection” to evaluate the synthetic data. Based on this metric, we filtered synthetic data from initially 69,443 pairs to 20,833 high quality pairs. Similar to our real-world dataset, our synthetic data is also capable of generating multiple reflection for a single scene, thereby enhancing the diversity and practicality of the training data.
Appendix D Implementation Details
We employ a progressive training strategy to divide the whole process into three distinct stages. All three stages are executed on 3 NVIDIA GeForce RTX 3090 GPUs with a batch size of 1. The pretrained weights of Stable Diffusion V2.1 are used to initialize the U-net and ControlNet. We employ the AdamW (Loshchilov and Hutter 2017) optimizer with a fixed learning rate of $3\times 10^{-4}$ except during the reflection-invariant finetuning phase, where a learning rate of $1\times 10^{-4}$ is used. To optimize memory usage during the reflection-invariant finetuning in the second phase, we alternate the training of the ControlNet and the upsampling blocks of the U-Net every 100 steps. The durations for the three parts are 2 days, 1 day, and 1 day, respectively. Our method achieves fast inference speed, processing a 768-resolution image in approximately 1 second on a single NVIDIA GeForce RTX 3090 GPU.
We evaluate the performance of our model on three established benchmarks: $Nature$ (Li et al. 2020), $Real$ (Zhang, Ng, and Chen 2018), and $SIR^{2}$ (Wan et al. 2017). Additionally, we introduce a new benchmark derived from our DRR dataset to provide a more comprehensive evaluation. Our DRR dataset includes one standard set and another challenging set, each containing 200 image pairs sampled from 20 different scenes. This design further validates the robustness and generalization capabilities of our approach across diverse real-world conditions.
Appendix E More Qualitative Comparison
We provide additional qualitative comparisons between our method and other approaches, as illustrated in Fig. 13 and Fig. 14, highlighting the superior performance of our comprehensive method.
Appendix F Dereflection Any Image
Appendix G Downstream application
@@
Methods Venue NatureNature (20) RealReal (20) $SIR^{2}$ (500) DRRDRR-$S$ (200) DRRDRR-$C$ (200)
[…]
RobustSIRR CVPR 2023 20.94 0.770 22.71 0.787 22.61 0.872 19.68 0.756 20.24 0.692
DSRNet ICCV 2023 24.86 0.823 23.31 0.791 25.65 0.919 22.33 0.846 21.93 0.820
RRW CVPR 2024 25.79 0.833 21.51 0.767 25.31 0.907 22.39 0.857 21.84 0.820
L-DiffER ECCV 2024 23.95 0.831 23.77 0.821 25.18 0.911 - - - -
DSIT NeurIPS 2024 26.25 0.833 24.54 0.814 26.34 0.922 23.48 0.869 22.46 0.817
RDNet CVPR 2025 25.94 0.835 25.10 0.833 26.38 0.924 23.31 0.865 22.25 0.823
RobustSIRR CVPR 2023 23.74 0.818 23.00 0.793 24.14 0.897 23.25 0.793 20.94 0.709
DSRNet ICCV 2023 25.29 0.827 24.52 0.822 25.97 0.925 24.47 0.874 22.13 0.826
RRW CVPR 2024 26.76 0.847 24.24 0.809 25.32 0.916 23.38 0.869 22.24 0.823
DSIT NeurIPS 2024 26.30 0.833 25.18 0.822 26.20 0.927 26.46 0.891 23.34 0.834
RDNet CVPR 2025 26.72 0.841 24.80 0.827 26.09 0.923 24.96 0.879 22.17 0.827
Ours - 27.05 0.846 25.24 0.840 27.32 0.931 27.21 0.901 23.83 0.840
Appendix H Limitations
As reflection removal is an ill-posed problem, failure cases still exists. When the reflection and target scene are seamlessly integrated, it is hard to distinguish which one is the target to be preserved. An example is shown in Fig. 12. To address this limitation, incorporating semantic or user guidance offers a viable solution and a promising direction for future research.
Figure 12: An example of limitation.
Figure 13: More qualitative comparison of our method with other approaches (not finetuned on our data setting), tested on $SIR^{2}$ (Wan et al. 2017).
Figure 14: Qualitative comparison of our method with other approaches (not finetuned on our data setting), tested on DRRDRR-$S$ and DRRDRR-$C$.